### v1

# Beer Path Problems in Temporal Graphs

Abstract
Computing paths in graph structures is a fundamental operation in a wide range of applications, from transportation networks to data analysis. The beer path problem, which captures the option of visiting points of interest, such as gas stations or convenience stops, prior to reaching the final destination, has been recently introduced and extensively studied in static graphs. However, existing approaches do not account for temporal information, which is often crucial in real-world scenarios. For instance, transit services may follow fixed schedules, and shops may only be accessible during certain hours.
In this work, we introduce the notion of beer paths in temporal graphs, where edges are time-dependent and certain vertices (beer vertices) are active only at specific time instances. We formally define the problems of computing earliest-arrival, latest-departure, fastest, and shortest temporal beer paths and propose efficient algorithms for these problems under both edge stream and adjacency list representations. The time complexity of each of our algorithms is aligned with that of corresponding temporal pathfinding algorithms, thus preserving efficiency.
Additionally, we present preprocessing techniques that enable efficient query answering under dynamic conditions, for example new openings or closings of shops. We achieve this through appropriate precomputation of selected paths or by transforming a temporal graph into an equivalent static graph.
keywords:
Beer path, Shortest paths, Dynamic graph queries, Temporal graphs, Graph algorithms
category:
[…]
1 Introduction
Answering path queries is a fundamental operation on graphs, with numerous optimization and information retrieval tasks relying on an efficient response to such queries. Bacic et al. [1] recently introduced the concepts of beer vertices and beer paths to model the option of taking a detour towards a set of points of interest when travelling from a source point to a destination point. In fact, choosing detours is an essential part of daily life and modern computing systems. When travelling from point A to point B, we often need to adjust our route, perhaps to refuel at a gas station or pick up a beer. Given a weighted graph $G=(V,E,w)$ and a subset $B\subseteq V$ of beer vertices, a beer path between two vertices $u$ and $v$ is a path that starts from $u$, ends at $v$, and visits at least one beer vertex in $B$. The beer distance between two vertices is the shortest length of any beer path connecting them.
Since their introduction, beer paths and distances have been extensively studied in different graph classes and frameworks [1, 2, 4, 5, 6, 7]. All existing techniques, however, are applicable to static graphs, which limits the uses of beer paths. In fact, because of the inherent nature of beer path applications, temporal information is a fundamental feature to consider in such scenarios. For example, bus routes may operate only at specific times, while beer shops may be subject to opening and closing hours. Temporal graphs [9, 10, 13, 21] extend the concept of static graphs to include such temporal information, meaning that each edge is associated with a starting time and a travelling time, and two vertices are adjacent only in the time instants in which an edge incident to them exists. In this work we introduce and study beer path problems in temporal graphs.
1.1 Related work
Temporal graphs have been the focus of research in several fields due to their ability to model a wide range of real-world scenarios. Comprehensive overviews on temporal graphs are provided in [8, 19]. The concepts of earliest arrival, latest departure, fastest, and shortest paths in temporal graphs were introduced in [20]. In the same paper, it is shown that all corresponding problems can be solved in time polynomial in the size of the temporal graph.
The problem of computing shortest beer paths in static graphs was recently introduced by Bacic et al. [1]. Since the problem is polynomially solvable, researchers have focused on specific graph classes to design data structures that allow answering beer path queries faster than the baseline method. Bacic et al. developed an index for outerplanar graphs [1]; Das et al. [5] studied the beer path problem on interval graphs; Hanaka et al. [7] achieved optimal query time on series-parallel graphs and linear preprocessing time on graphs having bounded-size triconnected components; Gudmundsson and Sha [6] focused on bounded treewidth graphs. Additionally, Bilò et al. [2] address the problem of constructing graph spanners w.r.t. the group Steiner metric, which generalizes the beer distance metric. The shortest beer path problem is a special case of the Generalized Shortest Paths problem and the NP-hard Generalized Travelling Salesperson Problem, both of which asking, for a given pair of vertices, to determine a path of minimum total cost that visits at least one location from each of a set of specified location categories (e.g., gas stations, grocery stores) [11, 15, 16, 17, 18].
1.2 Our contribution
In this paper, we introduce the concept of beer paths in temporal graphs and several related optimization problems, namely earliest-arrival, latest-departure, fastest, and shortest temporal beer paths (in short EABP, LDBP, FBP, SBP). In order to further incorporate temporal information that more accurately reflects real-life applications, our framework also considers timestamped beer vertices, i.e., the case where each beer vertex is augmented with the time instants [^1] We consider time instants and not intervals. With slight modifications, our algorithms can work for intervals as well, with logarithmic dependence on the number of intervals. in which it is active as a point of interest. This feature naturally models many scenarios, e.g., opening and closing times of bars or gas stations; active motorway toll booths; instants in which airports may serve for emergency landing, and so on.
In Section 3, we propose efficient algorithms for the aforementioned objectives, with each one differing both in methods and in time complexity. We define multiple-source earliest-arrival paths (MSEAP) and multiple-target latest-departure paths (MTLDP) and propose algorithms for calculating them, as an intermediate step for EABP and LDBP. [^2] MSEAPs and MTLDPs have interesting applications in their own right; for example, a MSEAP is the optimal path for sending an ambulance to some target destination, assuming an ambulance can start from multiple sources (hospitals), at a different time for each source. Using MSEAP subroutines, we propose two algorithms for EABP, one for the edge stream model and one for the adjacency list model, with the second one being significantly faster for graphs with no dominated edges. Similarly, we propose two algorithms for LDBP based on MTLDP subroutines. For FBP and SBP we rely on the computation of certain non-dominated paths via various domination criteria. For SBP in particular, we define a new path domination criterion and propose an algorithm for computing respective paths. The time complexity of our algorithms generally aligns with that of respective temporal path algorithms.
In Section 4, we propose preprocessing methods for answering beer path queries faster. For EABP and LDBP, this is achieved through the precomputation of selected non-dominated paths in the graph. For FBP and SBP, we utilize the graph transformation proposed by [20]. For FBP, this offers a speedup compared to the algorithm of Section 3. For SBP, no speedup is achieved, but we obtain a one-to-all SBP algorithm, which is significant because the respective algorithm of Section 3 only runs for a specific source and target.
2 Preliminaries and problem definition
2.1 Preliminaries
Let $G=(V,E)$ be a temporal graph, with $|V|=n$ and $|E|=M$. Each temporal edge $e\in E$ is defined as a quadruple $(u,v,t,\lambda)$, with $u,v\in V$, where $t$ denotes the starting time, $\lambda$ denotes the traversal time needed to travel from $u$ to $v$ starting at time $t$, and $t+\lambda$ represents the ending time. Most of our algorithms receive as input an edge stream, containing all edges sorted by ascending $t$-value. This is standard for temporal graphs [8, 12, 14, 20].
The set of temporal edges from $u$ to $v$ $(u,v\in V)$ is denoted by $E(u,v)$. Let $\pi=\max\limits_{u,v\in V}\{|E(u,v)|\}$ (maximum number of parallel edges) and $m=|\{u,v\in V:|E(u,v)|>0\}$ (number of connected pairs of vertices). We denote with $N_{out}(u)=\{v:(u,v,t,\lambda)\in E\}$ the set of out-neighbours of $u\in V$, and with $d_{out}$ the maximum out-degree in $G$. We similarly define $N_{in}(u)$ and $d_{in}$. Finally, $d_{max}$ denotes the maximum between $d_{out}$ and $d_{in}$.
A temporal path $P$ in a temporal graph $G$ is a sequence of vertices $P=\langle v_{1},v_{2},\dots,v_{\ell}\rangle$ where $(v_{i},v_{i+1},t_{i},\lambda_{i})\in E$ is the $i$-th temporal edge on $P$ for $1\leq i\leq\ell$, and $(t_{i}+\lambda_{i})\leq t_{i+1}$ for $1\leq i<\ell$. The quantity $(t_{\ell}+\lambda_{\ell})$ is known as the ending time of $P$, denoted by $end(P)$, while $t_{1}$ is the starting time of $P$, $start(P)$. The duration of $P$ is defined as $dura(P)=end(P)-start(P)$. Finally, the sum of the traversal time of each temporal edge of $P$ is known as the distance of the path, denoted as $dist(P)$.
Definition 2.1 (Dominated edge).
We say that a temporal edge $e=(u,v,t,\lambda)$ is dominated if there exists another edge $e^{\prime}=(u,v,t^{\prime},\lambda^{\prime})$ such that $t^{\prime}\geq t$ and $t^{\prime}+\lambda^{\prime}\leq t+\lambda$, with at least one of the two inequalities being strict. Symmetrically, we say that edge $e^{\prime}$ dominates edge $e$.
Definition 2.2 (Dominated paths).
Let $P$ be a temporal $u-v$ path. We say that $P$ is dominated if there exists another temporal $u-v$ path $P^{\prime}$ such that $\text{start}(P^{\prime})\geq\text{start}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict. We say that $P$ is distance-wise dominated if there exists a temporal $u-v$ path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
In some of our algorithms we will use a subroutine that computes all [^3] If there are paths $P,P^{\prime}$ with $start(P)=start(P^{\prime})$ and $end(P)=end(P^{\prime})$, the algorithm only stores one of them, since they can be used interchangeably for any objective. Throughout the paper, when we say ”all non-dominated paths”, we imply that only one path is kept when there are such ties. non-dominated paths from one node to all others. This is useful for finding fastest paths, as well as for precomputing certain paths in temporal graphs to answer queries faster. This subroutine consists of part of the fastest-path algorithm of [20]. We defer the full algorithm to Appendix A.
Algorithm 1 Non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
Algorithm 1 runs in time $\mathcal{O}(n+M\log(c))$, where $c=\min\{d_{in},d_{out}\}$. Note that $c$ is a bound for the amount of non-dominated pairs $(s,a)$ (start, arrival) between any two vertices and, thus, also a bound for the size of each $L_{v}$. For more details, we defer the reader to [20].
We will also use a similar subroutine for distance-wise non-dominated paths, which is part of the shortest-path algorithm of [20]. Its full description is deferred to Appendix A.
Algorithm 2 Dist_non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
Algorithm 2 runs in time $\mathcal{O}(n+M\log(d_{in}))$. Observe that $d_{in}$ is a bound for the amount of distance-wise non-dominated pairs $(d,a)$ (distance, arrival) between any two vertices.
2.2 The temporal beer path problem
We assume a temporal graph with a set $B=\{b_{1},\dots,b_{k}\}\subseteq V$ of $k$ beer vertices. For each beer vertex $b\in B$, we are given a set $T_{b}$ of timestamps denoting the time instants in which $b$ is active. We assume that beer vertices can still be traversed in time units in which they are inactive, functioning as regular vertices. We define $T=\max\limits_{b\in B}\{|T_{b}|\}$.
Given a temporal graph $G$, vertices $x,y\in V$ and a time interval $[t_{\alpha},t_{\omega}]$, we denote the set of temporal beer paths between two vertices $x$ and $y$ starting after $t_{\alpha}$ and ending before $t_{\omega}$ as $P_{B}(x,y,[t_{\alpha},t_{\omega}])=$$\{P=\langle x=v_{1},v_{2},\dots,v_{\ell}=y\rangle:P$ is a temporal path from $x$ to $y$ such that $start(P)\geq t_{\alpha},\ end(P)\leq t_{\omega},\ \exists\ (i\in[\ell],t\in T_{v_{i}}):(v_{i}\in B,\ t\in[t_{i}+\lambda_{i},t_{i+1}])\}$. Note that a beer path must traverse at least one beer vertex at a time in which it is active.
Definition 2.3 (EABP).
A temporal $x$-$y$ path $P\in{P_{B}}(x,y,[t_{\alpha},t_{\omega}])$ is an earliest-arrival beer path if $end(P)=\min\{end(P^{\prime}):P^{\prime}\in P_{B}(x,y,[t_{\alpha},t_{\omega}])\}$.
Definition 2.4 (LDBP).
A temporal $x$-$y$ path $P\in{P_{B}}(x,y,[t_{\alpha},t_{\omega}])$ is a latest-departure beer path if $start(P)=\max\{start(P^{\prime}):P^{\prime}\in P_{B}(x,y,[t_{\alpha},t_{\omega}])\}$.
Definition 2.5 (FBP).
A temporal $x$-$y$ path $P\in{P_{B}}(x,y,[t_{\alpha},t_{\omega}])$ is a fastest beer path if $dura(P)=\min\{dura(P^{\prime}):P^{\prime}\in P_{B}(x,y,[t_{\alpha},t_{\omega}])\}$.
Definition 2.6 (SBP).
A temporal $x$-$y$ path $P\in{P_{B}}(x,y,[t_{\alpha},t_{\omega}])$ is a shortest beer path if $dist(P)=\min\{dist(P^{\prime}):P^{\prime}\in P_{B}(x,y,[t_{\alpha},t_{\omega}])\}$.
Figure 1: Example in which the $x-y$ EABP/LDBP/FBP/SBP use $b_{1},\ldots,b_{4}$ respectively. Numbers in parentheses denote $(t,\lambda)$ of edges and numbers in brackets denote beer vertex active times. $x-b_{5}-y$ is not a valid temporal beer path because $4\notin T_{b_{5}}$. The $x-b_{2}$ edge with $(t,\lambda)=(18,9)$ is dominated and can be removed from the graph without affecting any objective.
3 Algorithms for temporal beer path problems
In all our algorithms, we focus on the computation of the objective function (e.g., the earliest arrival time) instead of reporting the corresponding optimal path. This path can be easily recovered by storing parent information.
3.1 Algorithms for earliest-arrival beer path
In this subsection we propose an EABP algorithm (from source $x$ to all nodes), based on the earliest-arrival time algorithm of [20] with a modification for calculating multiple-source earliest-arrival paths (MSEAP - Def. 3.1). Our algorithm makes two sequential passes of the graph’s edge stream, calculating the earliest arrival times from $x$ to each beer vertex in the first pass. For the second pass, all arrival times are initialized to $\infty$, except those of the beer vertices, which are initialized to the minimum active time greater than or equal to their respective earliest arrival time. We will later prove that this correctly calculates EABP times.
We also present a MSEAP subroutine for graphs containing no dominated edges. [^4] This is a reasonable assumption for temporal graphs, since dominated edges are not useful for any of the commonly considered objectives. This algorithm is similar-in-spirit to Dijkstra’s algorithm and uses binary search for parallel edges, achieving a running time independent of $M$. Although faster than the previous one, it requires random access to the edges of the graph (through e.g. an adjacency list with temporal logs [3]), instead of scanning an edge stream in a specific order.
3.1.1 An earliest-arrival beer path algorithm in edge stream representation
Definition 3.1 (MSEAP).
Given a temporal graph $G=(V,E)$ and initial times $init[v]\in\mathbb{R}^{+}\cup\{\infty\}$, $v\in V$, we say that a temporal path $x-y$ starting from $x$ no earlier than $init[x]$ is a multiple-source earliest-arrival path to $y$ if its arrival time at $y$ is minimum among all temporal paths $v-y$, $v\in V$, starting from $v$ no earlier than $init[v]$. We call this arrival time the multiple-source earliest-arrival path time of $y$.
Intuitively, a MSEAP is a path that can reach a target node in the earliest possible time unit, starting from any node of the graph, with restrictions for the starting time of each node. If some $init[v]$ is $\infty$, we cannot start a path from $v$. We first show that a generalization of the earliest-arrival path algorithm of [20] can compute MSEAPs.
Algorithm 3 MSEAP$(G,[t_{\alpha},t_{\omega}],init)$
1:A temporal graph $G$ in its edge stream representation, a time window $[t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
2:The MSEAP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
3:$\tau[v]\leftarrow init[v]$ for all $v\in V$
4:for $e=(u,v,t,\lambda)$ in the edge stream do
5:     if $\tau[u]\leq t$ and $t+\lambda\leq t_{\omega}$ then
6:         $\tau[v]\leftarrow\min\{\tau[v],t+\lambda\}$
7:return $\tau[v]$ for each $v\in V$
Algorithm 3 scans each temporal edge once and thus runs in time $\mathcal{O}(n+M)$.
{observation}
With initialization $init[x]=t_{\alpha}$ for some $x\in V$ and $init[v]=\infty$ for $v\in V\setminus\{x\}$, Algorithm 3 coincides with the earliest-arrival time algorithm of [20] with source node $x$ and calculates single-source earliest arrival paths.
Lemma 3.2 (Wu et al. [20]).
For all nodes $x,y\in V$, if there exists some temporal $x-y$ path, then there exists some earliest-arrival $x-y$ path $\langle v_{0}=x,v_{1},v_{2},\ldots,v_{k}=y\rangle$ for which every prefix-subpath $\langle x,v_{1},v_{2},\ldots,v_{i}\rangle$ is an earliest-arrival $x-v_{i}$ path.
{observation}
In temporal graphs, the property described in Lemma 3.2 only holds for some $x-y$ earliest-arrival path, in contrast to classical graphs, where a similar property holds for all $x-y$ shortest paths.
The following lemma is analogous to Lemma 3.2 for MSEAPs.
Lemma 3.3.
For $G=(V,E)$ and $init[v]\in\mathbb{R}^{+}\cup\{\infty\}$, $v\in V$, it holds that for every $y\in V$ with non-infinite MSEAP time there exists some MSEAP $\langle v_{0},v_{1},\ldots,v_{p}=y\rangle$ to $y$ for which every prefix-subpath $\langle v_{0},v_{1},\ldots,v_{i}\rangle$ is a MSEAP to $v_{i}$.
Proof 3.4.
Let $P=\langle u_{0},u_{1},\dots,u_{p}=y\rangle$ be an arbitrary MSEAP to $y$. If the desired property does not hold for $P$, take the maximum $i\in[p]$ s.t. the prefix-subpath $R=\langle u_{0},\dots,u_{i}\rangle$ of $P$ is not a MSEAP to $u_{i}$. Let $Q=\langle w_{0},w_{1},\dots,w_{j}=u_{i}\rangle$ be a MSEAP to $u_{i}$. Replace $R$ with $Q$, keeping the suffix $\langle u_{i+1},\dots,u_{p}\rangle$ unchanged. Since $Q$ reaches $u_{i}$ no later than $R$ (by definition of MSEAP), the resulting path $P^{\prime}$ is a valid temporal path and the arrival times of nodes $u_{i+1},\ldots u_{p}$ through $P^{\prime}$ remain the same as they were in $P$ (and, thus, the desired property is preserved for those nodes). This process can be repeated until every prefix-subpath is a MSEAP.
Lemma 3.5.
Algorithm 3 computes the MSEAP times to each vertex, within time window $[t_{\alpha},t_{\omega}]$. If time $\infty$ is returned for some vertex, then that vertex is unreachable from all $v\in V$ (starting no earlier than $init[v]$) within $[t_{\alpha},t_{\omega}]$.
Proof 3.6.
For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time (within $[t_{\alpha},t_{\omega}]$). Since $\tau[v]$ is initialized to $init[v]$, any update done by the algorithm calculates the arrival time of a valid temporal path, starting from some node $v$ at time $init[v]$. Thus, at any step of the algorithm it holds that $\tau[v]\geq\tau^{*}[v]$, $\forall v\in V$.
Fix an arbitrary vertex $y$ with $\tau^{*}[y]<\infty$. By Lemma 3.3, there exists a MSEAP $P=\langle v_{0}=x,v_{1},\dots,v_{p}=y\rangle$ to $y$ s.t. every prefix-subpath $\langle v_{0},v_{1},v_{2},\ldots,v_{i}\rangle$ of $P$ is a MSEAP to $v_{i}$. Denote the $i$-th edge of $P$ by $e_{i}=(v_{i-1},v_{i},t_{i},\lambda_{i}),i=1,2,\dots,p$. Thus, it holds that $t_{i}\geq\tau^{*}[v_{i-1}]$ and $\tau^{*}[v_{i}]=t_{i}+\lambda_{i}$. We will show by induction on $i$ that Algorithm 3 computes $\tau[v_{i}]=\tau^{*}[v_{i}]$ after processing $e_{i}$.
Base $(i=0)$. By the prefix-subpath property of $P$, we have $\tau^{*}[x]=init[x]$. By the initialization of the algorithm, $\tau[x]\leq init[x]$. Thus, $\tau[x]=\tau^{*}[x]$.
Inductive step $(i\geq 1)$. Assume $\tau[v_{i-1}]=\tau^{*}[v_{i-1}]$ after $e_{i-1}$ is processed. When Algorithm 3 scans $e_{i}=(v_{i-1},v_{i},t_{i},\lambda_{i})$, it holds that $t_{i}\geq\tau[v_{i-1}]=\tau^{*}[v_{i-1}]$. Thus, $\tau[v_{i}]$ is updated to $t_{i}+\lambda_{i}=\tau^{*}[v_{i}]$ (unless it was already equal to that value). Since it always holds that $\tau[v]\geq\tau^{*}[v]$, $\forall v\in V$, $\tau[v_{i}]$ will never be updated again after this step.
From the above we infer that $\tau[v]=\infty$ can be returned for some $v\in V$ only if no MSEAP to $v$ exists. This implies that $v$ is unreachable from all $u\in V$ (starting no earlier than $init[u]$) within $[t_{\alpha},t_{\omega}]$.
We now present an EABP algorithm, which calls Algorithm 3 twice, calculating earliest-arrival path times in the first phase and EABP times in the second phase.
Algorithm 4 EABP$(G,x,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G$, source vertex $x$, time window $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The EABP time from $x$ to every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
3:$init[x]\leftarrow t_{\alpha},\,init[v]\leftarrow\infty$ for all $v\in V\setminus\{x\}$
4:$\tau[v]\leftarrow\texttt{MSEAP($G,[t_{\alpha},t_{\omega}],init$)}$, for $v\in V$
5:$init^{\prime}[v]\leftarrow\infty$ for all $v\in V$
6:for all $b\in B$ do
7:     $beer\_time\leftarrow\min\{time\in T_{b}\mid time\geq\tau[b]\}$triangleright $\infty$ if it does not exist
8:     if $beer\_time\leq t_{\omega}$ then
9:         $init^{\prime}[b]\leftarrow beer\_time$
10:$\tau^{\prime}[v]\leftarrow\texttt{MSEAP($G,[t_{\alpha},t_{\omega}],init^{\prime}$)}$, for $v\in V$
11:return $\tau^{\prime}[v]$ for each $v\in V$
Theorem 3.7.
Algorithm 4 computes the EABP times from $x$ to all $v\in V$ in time $\mathcal{O}(n+M+k\log(T))$.
Proof 3.8.
Algorithm 4 scans all edges twice. A binary search is done for each of the $k$ lists $T_{b}$. Hence, Algorithm 4 runs in time $\mathcal{O}(n+M+k\log(T))$.
By Observation 3.1.1, the first call to Algorithm 3 correctly computes the earliest-arrival path times from $x$ to the beer vertices (within $[t_{\alpha},t_{\omega}]$). Thus, the binary search on $T_{b}$ computes for each $b\in B$ the earliest departure time from $b$ s.t. $b$ is used as the beer vertex of a beer path from $x$ to any other node (within $[t_{\alpha},t_{\omega}]$). Denote this time as $t[b]$, $b\in B$. Observe that, for any node $v$, the EABP time from $x$ to $v$ is the earliest time in which $v$ is reachable from some $b\in B$, starting from $b$ no earlier than $t[b]$. This coincides with the MSEAP time to $v$ for $init[b]=t[b]$, $b\in B$, and $init[u]=\infty$, $u\in V\setminus B$. Hence, the second call to Algorithm 3 correctly computes the EABP times from $x$ to all nodes (within $[t_{\alpha},t_{\omega}]$), by Lemma 3.5.
3.1.2 A faster earliest-arrival beer path algorithm with adjacency list representation
We present an alternative algorithm to be used as the MSEAP subroutine of Algorithm 4, in the place of Algorithm 3. Although faster for graphs with no dominated edges, this algorithm requires an adjacency list instead of an edge stream. The algorithm is similar-in-spirit to Dijkstra’s algorithm, using a min-priority queue for the nodes and their arrival times.
Algorithm 5 MSEAP_alternative$(G,[t_{\alpha},t_{\omega}],init)$
1:A temporal graph $G$ in adjacency list representation, a time window $[t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
2:The MSEAP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
3:$\tau[v]\leftarrow init[v]$ for all $v\in V$
4:$PQ\leftarrow\emptyset$triangleright a min-priority queue
5:$PQ.enqueue(\tau[v],v)$ for all $v\in V$triangleright Add $v$ to priority queue with $\tau[v]$ priority
6:while $PQ\neq\emptyset$ do
7:     $(\tau[u],u)\leftarrow PQ.extractMin()$
8:     for $v\in N_{out}(u)$ do
9:         Let $(u,v,t,\lambda)\in E(u,v)$ be a temporal edge with minimum $t$ s.t. $t\geq\tau[u]$
10:         if $t+\lambda\leq t_{\omega}$ and $t+\lambda<\tau[v]$ then
11:              $\tau[v]\leftarrow t+\lambda$
12:              $PQ.decreaseKey(v,\tau[v])$
13:return $\tau[v]$ for each $v\in V$
Theorem 3.9.
For graphs with no dominated edges, Algorithm 4 computes the EABP times from $x$ to all $v\in V$ in time $\mathcal{O}(m\log(\pi)+n\log(n)+k\log(T))$, if Algorithm 5 is used as the MSEAP subroutine.
The proof of Theorem 3.9 is similar-in-spirit to the proofs of Lemma 3.5 and Theorem 3.7 and is deferred to Appendix B. The $n\log n$ term in the time complexity stems from the usage of a Fibonacci heap as min-priority queue. The $m\log\pi$ term stems from binary search on parallel edges, which is possible by the assumption that $G$ contains no dominated edges.
{observation}
For graphs with no dominated edges, Algorithm 5 is a faster alternative to Algorithm 3 under reasonable assumptions (albeit requiring an adjacency list as input instead of an edge stream). $M$ is bounded by $m\pi$ and may be arbitrarily larger than $m$, which renders $\mathcal{O}(m\log(\pi))$ a significant improvement over $\mathcal{O}(M)$.
3.2 Algorithms for latest-departure beer path
We propose two LDBP algorithms (from all nodes to some node $y$), based on the earliest-arrival algorithms presented in subsection 3.1. The main idea is to reverse the direction of time, starting from the target node $y$ as late as possible and traversing temporal edges in reverse, while storing the latest time in which we can depart from each node to reach $y$. Due to space constraints, we defer the full descriptions of these algorithms to Appendix A and we briefly mention the main differences compared to those of subsection 3.1.
Definition 3.10 (MTLDP).
Given a temporal graph $G=(V,E)$ and final times $fin[v]\in\mathbb{R}^{+}\cup\{-\infty\}$, $v\in V$, we say that a temporal $x-y$ path arriving at $y$ no later than $fin[y]$ is a multiple-target latest-departure path from $x$ if its starting time from $x$ is maximum among all temporal paths $x-v$, $v\in V$, arriving at $v$ no later than $fin[v]$. We call this starting time the multiple-target latest-departure path time of $x$.
The main algorithm (analogous to Algorithm 4) sets $fin[y]$ to $t_{\omega}$ and $fin[v]$ to $\infty$- ∞ for all other nodes. A MTLDP subroutine is used to calculate latest-departure times in the first phase and LDBP times in the second phase. A search in each list $T_{b}$ is done before the second phase, similar to Algorithm 4.
We now describe the main differences of the two MTLDP subroutines, compared to Algorithms 3 and 5. In the edge stream algorithm, the stream is scanned in reverse (i.e., by decreasing $t$-value) and each edge $(u,v,t,\lambda)$ updates $\tau[u]$ to $t$ if $\tau[u]<t$ and $t\geq t_{a}$ and $t+\lambda\leq\tau[v]$. In the adjacency list algorithm, a max-priority queue is used for MTLDP times and the binary search on parallel edges considers their $(t+\lambda)$-value [^5] Note that non-dominated edges sorted by $t$-value are also sorted by $(t+\lambda)$-value (by definition). Thus, this binary search is possible with no additional precomputation.. For each node $v$ extracted from the queue we consider its set $N_{in}(v)$ and the updates to the queue are done through increase-key operations. The time complexity of each algorithm is the same as in subsection 3.1 and the correctness proofs are analogous.
3.3 Algorithm for fastest beer path
A significant difference for FBP compared to EABP/LDBP is that no property similar to Lemma 3.2 holds. Indeed, computing a fastest path from the source to a beer vertex is unhelpful as an intermediate step for computing FBPs.
Lemma 3.11.
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is an $x-y$ FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
The proof of Lemma 3.11 is deferred to Appendix B. Based on Lemma 3.11, we propose a two-phase FBP algorithm (from source $x$ to all nodes), utilizing Algorithm 1. In the first phase, all non-dominated paths from $x$ to each $b\in B$ are calculated. As in subsection 3.1, we find the minimum eligible active time for the respective beer vertex for each of these paths. For the second phase, a dummy vertex is used as source, connected with each $b\in B$ by (multiple) dummy edges with starting and traversal times dictated by the values calculated previously for each path. Algorithm 1 is then used for the new graph, calculating eligible non-dominated paths from the dummy vertex to all vertices.
Algorithm 6 FBP$(G,x,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The FBP duration from $x$ to every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
3:$L_{b}\leftarrow\texttt{Non-dom\textunderscore paths($G,x,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright calculate non-dom paths with Alg 1
4:for $b\in B$ do
5:     for $(s,a)\in L_{b}$ do
6:         $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$triangleright $\infty$ if it does not exist
7:         if $a^{\prime}\leq t_{\omega}$ then
8:              Replace $(s,a)$ with $(s,a^{\prime})$ in $L_{b}$
9:         else
10:              Remove $(s,a)$ from $L_{b}$triangleright impossible to get beer
11:Add dummy node $\epsilon$ to the graph. Add an edge $(\epsilon,b,s,a-s)$ for each $(s,a)\in L_{b}$ for each $b\in B$; Let $G^{\prime}=(V^{\prime},E^{\prime})$ be the modified graph
12:$L_{v}^{\prime}\leftarrow\texttt{Non-dom\textunderscore paths($G^{\prime},\epsilon,[t_{\alpha},t_{\omega}]$)}$, for $v\in V$
13:for $v\in V$ do
14:     $f[v]\leftarrow s-a$, where $(s,a)\in L_{v}^{\prime}$ is the tuple with minimum difference
15:return $f[v]$ for each $v\in V$
Recall that Algorithm 1 runs in time $\mathcal{O}(n+M\log(c))$, where $c=\min\{d_{in},d_{out}\}$ is a bound for the size of each list $L_{v}$. The number of temporal edges added to the graph is bounded by $kc$, since there are $k$ lists $L_{b},\ b\in B$. Observe that $d_{in}$ can at most double itself after inserting the new edges, so the $c$-value of $G^{\prime}$ is bounded by $2d_{in}$. Thus, the second call to Algorithm 1 runs in time $\mathcal{O}(n+(M+kc)\log(d_{in}))$. Taking into account the binary searches for the active beer times, we obtain that Algorithm 6 runs in time $\mathcal{O}(n+(M+kc)\log(d_{in})+kc\log(T))$. [^6] The last for-loop of Alg. 6 can be incorporated into the second call of Alg. 1 without increasing its time complexity, by storing the tuples with minimum difference. We only present the pseudocode in this manner to help readability. The correctness of Algorithm 6 follows from Lemma 3.11.
3.4 Algorithm for shortest beer path
SBPs are significantly more complicated, in part because, somewhat unexpectedly, Lemma 3.11 does not hold verbatim for distance-wise non-dominated paths. To overcome this, we define the following alternative domination criterion.
Definition 3.12 (Inverse-distance-wise dominated path).
Let $P$ be a temporal $u-v$ path. We say that $P$ is inverse-distance-wise dominated if there exists another temporal $u-v$ path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{start}(P^{\prime})\geq\text{start}(P)$, with at least one of the inequalities being strict.
Lemma 3.13.
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is a $x-y$ SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t. $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$).
The proof of Lemma 3.13 is deferred to Appendix B. We propose a modification of Algorithm 2 to calculate inverse-distance-wise non-dominated paths from all vertices to some vertex $y$, using a reverse edge stream. Its full description is deferred to Appendix A.
Algorithm 7 Inv_dist_non-dom_paths($G,y,[t_{\alpha},t_{\omega}]$)
1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $[t_{\alpha},t_{\omega}]$.
Algorithm 7 runs in time $\mathcal{O}(n+M\log(d_{out}))$. Observe that $d_{out}$ is a bound for the amount of inverse-distance-wise non-dominated pairs $(d,s)$ (distance, start) between any two vertices and, thus, also a bound for the size of each list $I_{v}$.
Theorem 3.14.
Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $[t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
The proof of Theorem 3.14 is deferred to Appendix B. We propose a SBP algorithm, utilizing Algorithms 2 and 7. Unlike most of our algorithms, Algorithm 8 only calculates a path between two nodes; this happens because it uses two different domination criteria (see Lemma 3.13), with each one imposing different restrictions via its respective algorithm.
Algorithm 8 SBP$(G,x,y,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G=(V,E)$ in edge stream representation (both regular and reverse), source vertex $x$, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The SBP distance from $x$ to $y$ within $[t_{\alpha},t_{\omega}]$.
3:$dist\leftarrow\infty$
4:$L_{b}\leftarrow\texttt{Dist\_Non-dom\_paths($G,x,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright Alg 2 from $x$ to $b\in B$
5:$I_{b}\leftarrow\texttt{Inv\textunderscore dist\textunderscore Non-dom\textunderscore paths($G,y,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright Alg 7 from $b\in B$ to $y$
6:for $b\in B$ do
7:     for $(d,a)\in L_{b}$ do
8:         $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$triangleright $\infty$ if it does not exist
9:         if $a^{\prime}\leq t_{\omega}$ then
10:              Replace $(d,a)$ with $(d,a^{\prime})$ in $L_{b}$
11:         else
12:              Remove $(d,a)$ from $L_{b}$triangleright impossible to get beer
13:     for $(d,a)\in L_{b}$ do
14:         Find $(d^{\prime},s)$ in $I_{b}$ with minimum $s$ such that $s\geq a$triangleright find best fitting path
15:         $dist\leftarrow\min\{d+d^{\prime},dist\}$
16:return dist
Algorithm 8 calls Algorithms 2 and 7 once each. Recall that the size of each $L_{v}$ is bounded by $d_{in}$ and the size of each $I_{v}$ is bounded by $d_{out}$. Hence, Algorithm 8 runs in time $\mathcal{O}(n+M\log(d_{max})+kd_{in}(\log(T)+\log(d_{out})))$. The correctness of Algorithm 8 follows from Lemma 3.13 and Theorem 3.14.
4 Beer path queries with preprocessing
In this section we present preprocessing techniques that help improve the asymptotic complexity of answering beer path queries. This is useful for applications where multiple queries need to be answered, each with different source and target vertices, different time intervals and even different active beer shops. A natural application, for example, would be to quickly recompute a beer path assuming some beer shops just closed (or some new ones opened), without computing everything from scratch.
We assume that each query contains as input a source node $x$, a target node $y$, a time interval $[t_{\alpha},t_{\omega}]$ and a Boolean array $A=[a_{1},\ldots,a_{k}]$, with $a_{i}=\texttt{true}$ iff beer vertex $b_{i}$ is active. For simplicity, we assume that each beer vertex $b\in B$ is either active or inactive for the entire duration of a query. We also assume that the set $B$ containing the $k$ beer vertices is known during preprocessing; these vertices may be active or inactive for different queries, but no vertex $v\notin B$ may become a beer vertex.
4.1 EABP and LDBP with precomputation of non-dominated paths
We rely on the observation that Lemma 3.11 also holds for EABPs and LDBPs, with the same proof (although not for SBPs). This allows us to run Algorithm 1 to precompute non-dominated paths to and from beer vertices. To compute all non-dominated paths to each beer vertex (starting from any vertex) and all non-dominated paths from each beer vertex (to any vertex), it suffices to run Algorithm 1 $\mathcal{O}(n)$ times. Recall that the amount of non-dominated paths between any two vertices is bounded by $c=\min\{d_{in},d_{out}\}$. We infer that this preprocessing requires $\mathcal{O}(n^{2}+nM\log(c))$ time and $\mathcal{O}(nkc)$ space.
With this preprocessing, any EABP query can be answered in the following manner. We need a binary search for each $b_{i}\in B$ with $a_{i}=\texttt{true}$ to find the non-dominated $x-b_{i}$ path that starts as early as possible (but no earlier than $t_{\alpha}$). Then, for each $b_{i}\in B$ with $a_{i}=\texttt{true}$ we need a binary search to find the non-dominated $b_{i}-y$ path that starts as early as possible (but no earlier than the arrival time of the respective $x-b_{i}$ path). Concatenating these paths with the previous ones yields $k$ candidate EABPs, out of which we pick the best one. This process runs in $\mathcal{O}(k\log(c))$ time. Identical arguments hold for LDBP queries.
4.2 Preprocessing through graph transformation
Figure 2: Graph transformation, from $G$ in (a) to $\tilde{G}$ in (b).
In this subsection we adopt the graph transformation approach proposed in [20] as preprocessing. This technique converts a temporal graph $G=(V,E)$ into a static, weighted, directed acyclic graph $\tilde{G}=(\tilde{V},\tilde{E})$ that explicitly encodes temporal constraints. The key idea is to model the timing of events in the temporal graph by creating time-stamped copies of each node and connecting them with appropriately weighted edges. The transformed graph $\tilde{G}$ contains $\mathcal{O}(M)$ vertices and edges and can be produced from $G$ in $\mathcal{O}(M\log(d_{max}))$ time. For each $v\in V$, the transformation creates two sets of time-stamped nodes: $\tilde{V}_{in}(v)$, which includes a node $(v,t)$ for every unique time $t$ at which an edge arrives at $v$, and $\tilde{V}_{out}(v)$, which includes a node $(v,t)$ for every time $t$ at which an edge departs from $v$. These sets are sorted in descending time order. Figures 2(a),(b) show a temporal graph $G$ and its transformed graph $\tilde{G}$, respectively. We defer the reader to Appendix C for more details regarding this transformation.
4.2.1 Fastest beer path on transformed graph
Let $S=\{(x,t):(x,t)\in\tilde{V}_{out}(x),\ t_{\alpha}\leq t\leq t_{\omega}\}$ be the set of time-stamped copies of $x$, sorted in descending order of $t$. In order to calculate fastest temporal paths from $x$ to every $v\in V$, it suffices to start a BFS from each $(x,t)\in S$ (in descending order of $t$), skipping nodes already explored in previous BFSs to avoid redundancy [^7] Reaching $(v,t)$ from $(x,t_{1})$ results in smaller duration compared to reaching it from $(x,t_{2})$ with $t_{2}<t_{1}$.. The duration of each path is calculated as the difference between the timestamp of the target and the timestamp of the source [20]. We will modify this technique to compute FBPs on $\tilde{G}$.
For each vertex $(v,t)\in\tilde{G}$, we keep track of a $3$-valued flag, indicating whether $(v,t)$ has been visited by a beer path (value $2$) or a regular path (value $1$) or not at all (value 00), with higher values superseding smaller ones. We run BFSs from each $(x,t)\in S$ in descending order of $t$ (as described above), setting the flag of $(x,t)$ to $1$. We only visit a vertex if its flag is smaller than the parent’s flag; in this case, the child obtains the parent’s flag. If the child is a copy of some $b_{i}\in B$ with $a_{i}=\texttt{true}$, we always update its flag to $2$. When the flag of a vertex becomes $2$, we calculate the duration of the respective beer path as the difference between the timestamp of that vertex and the timestamp of the source.
The above answers FBP queries on $\tilde{G}$ in $\mathcal{O}(M)$ time, since each $v\in\tilde{G}$ is visited at most twice. The correctness of the algorithm follows from the fact that $\tilde{G}$ fully encodes all possible temporal walks in $G$, combined with arguments similar to the ones for regular fastest paths.
{observation}
EABP and LDBP queries can also be answered in $\tilde{G}$ in $\mathcal{O}(M)$ time with similar techniques. However, this is not particularly useful, since the algorithms proposed in subsections 3.1 and 3.2 require no preprocessing and are just as fast under reasonable assumptions (or even faster, depending on representation).
4.2.2 One-to-all shortest beer path on transformed graph
To compute shortest temporal paths from $x\in V$, it suffices to run Dijkstra’s algorithm in $\tilde{G}$ from a dummy vertex $x^{\prime}$, connected via directed edges of weight 0 to every $(x,t)\in\tilde{V}_{\text{out}}(x)$ s.t. $\ t_{\alpha}\leq t\leq t_{\omega}$, thus representing all valid departure times from $x$ [20].
We modify Dijkstra’s algorithm by storing a $3$-valued flag for each $(v,t)\in\tilde{G}$ and updating it in the same manner described in the previous subsection. This algorithm visits each $(v,t)\in\tilde{G}$ at most twice, thus calculating SBP distances in $\mathcal{O}(M\log(M))$ time.
{observation}
Although this SBP algorithm is not faster than Algorithm 8 (under reasonable assumptions), it computes SBPs from a source to all other nodes.
Appendix A Omitted Algorithms
A.1 Non-dominated path algorithms from Wu et al.
Algorithm 1 Non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
3:for $v\in V$ do
4:     Create a sorted list $L_{v}$, where an element of $L_{v}$ is a pair $(s,a)$ in which $s$ is the starting time of a path $P$ from $x$ to $v$ and is used as the key for ordering in $L_{v}$, and $a$ is the time in which the path $P$ arrives at $v$; initially, $L_{v}$ is empty
5:for $e=(u,v,t,\lambda)$ in the edge stream do
6:     if $t\geq t_{\alpha}$ and $t+\lambda\leq t_{\omega}$ then
7:         if $u=x$ and $(t,t)\not\in L_{x}$ then
8:              Insert $(t,t)$ into $L_{x}$
9:         Let $(s^{\prime},a^{\prime})$ be the element in $L_{u}$ for which $a^{\prime}=\max\{a\mid(s,a)\in L_{u},\ a\leq t\}$
10:         if $\exists(s,a)\in L_{v}$ such that $s=s^{\prime}$ and $a>t+\lambda$ then
11:              Replace $(s,a)$ with $(s^{\prime},t+\lambda)$
12:         else
13:              Insert $(s^{\prime},t+\lambda)$ into $L_{v}$
14:         Remove dominated elements in $L_{v}$
15:return $L_{v}$ for all $v\in V$
Algorithm 2 Dist_non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
3:for $v\in V$ do
4:     Create a sorted list $L_{v}$, where an element of $L_{v}$ is a pair $(d,a)$ in which $d$ is the distance of a path $P$ from $x$ to $v$ and $a$ is the time in which path $P$ arrives at $v$ and is used as the key for ordering in $L_{v}$; initially, $L_{v}$ is empty;
5:for $e=(u,v,t,\lambda)$ in the edge stream do
6:     if $t\geq t_{\alpha}$ and $t+\lambda\leq t_{\omega}$ then
7:         if $u=x$ and $(0,t)\not\in L_{x}$ then
8:              Insert $(0,t)$ into $L_{x}$
9:         Let $(d^{\prime},a^{\prime})$ be the element in $L_{u}$ for which $a^{\prime}=\max\{a\mid(d,a)\in L_{u},\ a\leq t\}$
10:         if $\exists(d,a)\in L_{v}$ such that $a=t+\lambda$ and $d>d^{\prime}+\lambda$ then
11:              Replace $(d,a)$ with $(d^{\prime}+\lambda,t+\lambda)$
12:         else
13:              Insert $(d^{\prime}+\lambda,t+\lambda)$ into $L_{v}$
14:         Remove distance-wise dominated elements in $L_{v}$
15:return $L_{v}$ for all $v\in V$
A.2 Modifed non-dominated path algorithm for new domination criterion
Algorithm 7 Inv_dist_non-dom_paths($G,y,[t_{\alpha},t_{\omega}]$)
1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $[t_{\alpha},t_{\omega}]$.
3:for $v\in V$ do
4:     Create a sorted list $I_{v}$, where an element of $I_{v}$ is a pair $(d,s)$ in which $d$ is the distance of a path $P$ from $v$ to $y$ and $s$ is the time in which path $P$ starts from $v$ and is used as the key for ordering in $I_{v}$; initially, $I_{v}$ is empty;
5:for $e=(u,v,t,\lambda)$ in the reverse edge stream do
6:     if $t\geq t_{\alpha}$ and $t+\lambda\leq t_{\omega}$ then
7:         if $v=y$ and $(0,t+\lambda)\not\in I_{y}$ then
8:              Insert $(0,t+\lambda)$ into $I_{y}$
9:         Let $(d^{\prime},s^{\prime})$ be the element in $I_{v}$ such that $s^{\prime}=\min\{s\mid(d,s)\in I_{v},\ s\geq t+\lambda\}$
10:         if $\exists\,(d,s)\in I_{u}$ such that $s=t$ and $d>d^{\prime}+\lambda$ then
11:              Replace $(d,s)$ with $(d^{\prime}+\lambda,t)$
12:         else
13:              Insert $(d^{\prime}+\lambda,t)$ into $I_{u}$
14:         Remove inverse-distance-wise dominated elements in $I_{u}$
15:return $I_{v}$ for all $v\in V$
A.3 Omitted MTLDP and latest-departure beer path algorithms
Algorithm 9 MTLDP$(G,[t_{\alpha},t_{\omega}],fin)$
1:A temporal graph $G$ in its reverse edge stream representation, a time window $[t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=-\infty$, for all $v\in V$.
2:The MTLDP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
3:$\tau[v]\leftarrow fin[v]$ for all $v\in V$
4:for $e=(u,v,t,\lambda)$ in the reverse edge stream do
5:     if $t\geq t_{\alpha}$ and $t+\lambda\leq\tau[v]$ then
6:         $\tau[u]\leftarrow\max\{\tau[u],t\}$
7:return $\tau[v]$ for each $v\in V$
Algorithm 10 MTLDP_alternative$(G,[t_{\alpha},t_{\omega}],fin)$
1:A temporal graph $G$ in adjacency list representation, a time window $[t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=-\infty$, for all $v\in V$.
2:The MTLDP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
3:$\tau[v]\leftarrow fin[v]$ for all $v\in V$
4:$PQ\leftarrow\emptyset$triangleright a max-priority queue
5:$PQ.enqueue(\tau[v],v)$ for all $v\in V$triangleright Add $v$ to priority queue with $\tau[v]$ priority
6:while $PQ\neq\emptyset$ do
7:     $(\tau[v],v)\leftarrow PQ.extractMax()$
8:     for $u\in N_{in}(v)$ do
9:         Let $(u,v,t,\lambda)\in E(u,v)$ be a temporal edge with maximum $t$ s.t. $t+\lambda\leq\tau[v]$triangleright binary search on $t+\lambda$ of $e\in E(u,v)$
10:         if $t\geq t_{\alpha}$ and $t>\tau[u]$ then
11:              $\tau[u]\leftarrow t$
12:              $PQ.increaseKey(u,\tau[u])$
13:return $\tau[v]$ for each $v\in V$
Algorithm 11 LDBP$(G,x,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G$, source vertex $x$, a time window $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The LDBP time to $y$ from every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
3:$fin[y]\leftarrow t_{\omega},\,init[v]\leftarrow-\infty$ for all $v\in V\setminus\{x\}$
4:$\tau[v]\leftarrow\texttt{MTLDP($G,[t_{\alpha},t_{\omega}],fin$)}$, for $v\in V$
5:$fin^{\prime}[v]\leftarrow-\infty$ for all $v\in V$
6:for all $b\in B$ do
7:     $beer\_time\leftarrow\max\{time\in T_{b}\mid time\leq\tau[b]\}$triangleright $\infty$- ∞ if it does not exist
8:     if $beer\_time\geq t_{\alpha}$ then
9:         $fin^{\prime}[b]\leftarrow beer\_time$
10:$\tau^{\prime}[v]\leftarrow\texttt{MTLDP($G,[t_{\alpha},t_{\omega}],fin^{\prime}$)}$, for $v\in V$
11:return $\tau^{\prime}[v]$ for each $v\in V$
Appendix B Omitted proofs
Theorem B.14.
For graphs with no dominated edges, Algorithm 4 computes the earliest-arrival beer path times from $x$ to all $v\in V$ in time $\mathcal{O}(m\log(\pi)+n\log(n)+k\log(T))$, if Algorithm 5 is used as the MSEAP subroutine.
Proof B.15.
Algorithm 5 uses a binary search [^8] Binary search is possible due to the assumption that there are no dominated edges in the graph. on each non-empty set $E(u,v)$ $(u,v\in V)$; this requires total time $\mathcal{O}(m\log(\pi))$. If a Fibonacci heap is used as priority queue, $\mathcal{O}(\log(n))$ time is required for each extract-min operation, while decrease-key operations are done in constant time. Thus, Algorithm 5 runs in time $\mathcal{O}(m\log(\pi)+n\log(n))$. Recall that Algorithm 4 calls the MSEAP subroutine twice, while also doing $\mathcal{O}(k)$ binary searches on lists of size $\mathcal{O}(T)$. We infer that Algorithm 4 runs in time $\mathcal{O}(m\log(\pi)+n\log(n)+k\log(T))$, if Algorithm 5 is used as the MSEAP subroutine.
We now prove the correctness of Algorithm 5 by induction. We only describe the induction briefly, since it is similar to the proof of correctness of Dijkstra’s classic algorithm. For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time within $[t_{\alpha},t_{\omega}]$. By definition of MSEAP, for the node with minimum $init[u]$ it is $\tau^{*}[u]=init[u]$; thus, the first node extracted from the queue has $\tau[u]=\tau^{*}[u]$. For each $v\in N_{out}(u)$, $\tau[v]$ is then updated through the $u-v$ edge that arrives at $v$ at the earliest possible time. Assuming the graph has no dominated edges, this is an edge with minimum $t$ s.t. $t\geq\tau[u]$. Inductively, any $u\in V$ extracted from the queue has $\tau[u]=\tau^{*}[u]$. This happens because paths that only use processed nodes have been considered in previous steps and it is impossible to arrive at $u$ earlier than $\tau[u]$ by traversing an unprocessed node $v\in V$ with $\tau[v]\geq\tau[u]$. Regarding the update of $\tau[v]$ for $v\in N_{out}(u)$, the same things hold as in the base of the induction, thus completing the proof.
The correctness of Algorithm 4 using Algorithm 5 as the MSEAP subroutine is then identical to the proof of Theorem 3.7.
Lemma B.16.
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is an $x-y$ FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
Proof B.17.
Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b,\ldots,u_{p^{\prime}}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$. Define the subpaths of $P$: $Q=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b\rangle$, $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\ldots,u_{p^{\prime}}=y\rangle$. If $Q$ is a dominated path, then there is some other path $Q^{\prime}=\langle w_{0}=x,w_{1},\ldots,w_{f}=b\rangle$ such that $start(Q^{\prime})\geq start(Q)$ and $end(Q^{\prime})\leq end(Q)$, with at least one of the inequalities being strict. Concatenate $Q^{\prime}$ with $R$ to obtain a $x-y$ path $P^{\prime}$. By the first inequality we obtain $dura(P^{\prime})\leq dura(P)$. By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and that $P^{\prime}$ is a valid temporal path. Thus, $P^{\prime}$ is a beer path with smaller or equal duration than $P$ and is also within $[t_{\alpha},t_{\omega}]$. Applying a similar argument for $R$ completes the proof.
Lemma B.18.
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is a $x-y$ SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t. $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$).
Proof B.19.
Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b,\dots,u_{p^{\prime}}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at some time $t\in T_{b}$. Define the sub-paths of $P$: $Q=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b\rangle$ and $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\dots,u_{p^{\prime}}=y\rangle$. If $Q$ is distance-wise dominated (within $[t_{\alpha},t_{\omega}]$), there exists a path $Q^{\prime}=\langle w_{0}=x,w_{1},\dots,w_{f}=b\rangle$ (within $[t_{\alpha},t_{\omega}]$), such that $\text{dist}(Q^{\prime})\leq\text{dist}(Q)$ and $\text{end}(Q^{\prime})\leq\text{end}(Q)$, with at least one of the two inequalities strict. Concatenate $Q^{\prime}$ with $R$ to obtain a path $P^{\prime}$ from $x$ to $y$. Because of the first inequality we get $\text{dist}(P^{\prime})\leq\text{dist}(P)$. By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and is a valid temporal path. Hence, $P^{\prime}$ is a beer path within $[t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P$.
If $R$ is inverse-distance-wise dominated (within $[t_{\alpha},t_{\omega}]$), there is a path $R^{\prime}=\langle z_{0}=b,z_{1},\dots,z_{g}=y\rangle$ (within $[t_{\alpha},t_{\omega}]$) such that $\text{dist}(R^{\prime})\leq\text{dist}(R)$ and $\text{start}(R^{\prime})\geq\text{start}(R)$, again with at least one inequality being strict. Concatenate $Q^{\prime}$ with $R^{\prime}$ to obtain a $x-y$ path $P^{\prime\prime}$. By the first inequality we obtain $\text{dist}(P^{\prime\prime})\leq\text{dist}(P^{\prime})$. By the second inequality, we obtain that $P^{\prime\prime}$ is a valid temporal beer path. Therefore, $P^{\prime\prime}$ is a beer path within $[t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P^{\prime}$.
Theorem B.20.
Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $[t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
Proof B.21.
Let $P=\langle v_{0}=v,v_{1},\dots,v_{p}=y\rangle$ be an inverse-distance-wise non-dominated path from $v$ to $y$ within $[t_{\alpha},t_{\omega}]$. For $i\in\{0,\ldots,p-1\}$, let $s_{i}$ be the time in which $P$ departs from $v_{i}$ and $d_{i}$ be the distance between $v_{i}$ and $y$ in $P$. We will prove that $(d_{0},s_{0})$ is contained in the list $I_{v}$ returned by Algorithm 7. Since we use a reverse edge stream, the edges of $P$ are processed in reverse order (starting from $y$). We use induction on the edges of $P$.
Base. Let $e_{p}=(v_{p-1},y,t_{p},\lambda_{p})$ be the last edge of $P$. When $e_{p}$ is processed, the algorithm inserts $(0,t_{p}+\lambda_{p})$ to $I_{y}$. Then, $(\lambda_{p},t_{p})$ will be inserted to $I_{v_{p-1}}$ unless that list already contains a pair that dominates $(\lambda_{p},t_{p})$ or one that is identical to it. In either case, $I_{v_{p-1}}$ will end up containing an element $(d,s)$ with $d\leq\lambda_{p}=d_{p-1}$ and $s\geq t_{p}=s_{p-1}$. Additionally, $I_{v_{p-1}}$ contains no dominated elements (because the algorithm removes them at the end of each loop).
Inductive step. Assume that $I_{v_{i}}$ has no dominated elements and contains an element $(d,s)$ with $d\leq d_{i}$ and $s\geq s_{i}$ (induction hypothesis). Let $e_{i}=(v_{i-1},v_{i},t_{i},\lambda_{i})$ be the $i$-th edge of $P$ and note that $s_{i}\geq t_{i}+\lambda_{i}$ because $P$ is a valid temporal path. When $e_{i}$ is processed, the algorithm finds the best fitting pair in $I_{v_{i}}$, i.e. $(d^{*},s^{*})$ such that $s^{*}$ is the minimum possible start larger than $t_{i}+\lambda_{i}$. By the induction hypothesis, such an $s^{*}$ exists and $d^{*}\leq d_{i}$. Thus, after $e_{i}$ is processed, $I_{v_{i-1}}$ will contain an element $(d,s)$ such that $d\leq d_{i}+\lambda_{i}=d_{i-1}$ and $s\geq t_{i}=s_{i-1}$. Also, $I_{v_{i-1}}$ will contain no dominated elements.
From the above induction we infer that $I_{v}$ contains an element $(d,s)$ with $d\leq d_{0}$ and $s\geq s_{0}$. By assumption, $P$ is an inverse-distance-wise non-dominated path within $[t_{\alpha},t_{\omega}]$, which implies that neither of the above inequalities can be strict. Hence, $I_{v}$ contains $(d_{0},s_{0})$, which finishes the correctness proof.
For the running time of the algorithm, observe that the amount of inverse-distance-wise non-dominated pairs from $x\in V$ to $y\in V$ is bounded by the out-degree of $x$. This implies that the size of any list $I_{v}$ is bounded by $d_{out}$ and, thus, search/update operations in $I_{v}$ take $\mathcal{O}(\log(d_{out}))$ time. Note that for each list $I_{v}$, at most one element is added for each of the outgoing edges of $v$. Thus, the total time of removing dominated elements from $I_{v}$ is bounded by the out-degree of $v$. For all lists $I_{v},\ v\in V$, this adds up to $\mathcal{O}(M)$ time. From all the above combined with the fact that Algorithm 7 processes each temporal edge once, we obtain that it runs in $\mathcal{O}(n+M\log(d_{out}))$ time.
Appendix C The graph transformation of Wu et al.
In this section we describe in detail the graph transformation of [20], which converts a temporal graph $G=(V,E)$ into an equivalent static, weighted, directed acyclic graph $\tilde{G}=(\tilde{V},\tilde{E})$.
For each vertex $v\in V$, the transformation creates two sets of time-stamped nodes: $\tilde{V}_{in}(v)$, which includes a node $(v,t)$ for every unique time $t$ at which an edge arrives at $v$, and $\tilde{V}_{out}(v)$, which includes a node $(v,t)$ for every time $t$ at which an edge departs from $v$. These sets are sorted in descending time order. Edges are then created in three categories. First, for each node in $\tilde{V}_{\text{in}}(u)$, a zero-weight edge is created to the earliest node $(v,t_{out})$ in $\tilde{V}_{out}(v)$ that has a time greater than or equal to it, if no edge from any other $(u,t^{\prime}_{\text{in}})\in\tilde{V}_{in}(u)$ to $(v,t_{out})$ has been created. Second, zero-weight edges are added between consecutive time-stamped nodes within $\tilde{V}_{\text{in}}(v)$ and within $\tilde{V}_{\text{out}}(v)$ to allow temporal “waiting” at the same vertex. Finally, for each original temporal edge $(u,v,t,\lambda)\in E$, a directed edge with weight $\lambda$ is added from $(u,t)\in\tilde{V}_{\text{out}}(u)$ to $(v,t+\lambda)\in\tilde{V}_{\text{in}}(v)$. Therefore, the graph $\tilde{G}$ has a size complexity of $\mathcal{O}(M)$ in both the number of vertices and edges.
The complexity of the transformation from the temporal graph $G=(V,E)$ to the static weighted graph $\tilde{G}$ primarily depends on two steps: vertex creation and edge creation. During vertex creation, for each vertex $v\in V$, time-stamped copies are created corresponding to distinct arrival and departure times. Sorting these time instances for all vertices takes $\mathcal{O}(M\log(d_{\max}))$ time, where $d_{\max}$ is the maximum degree (either in-degree or out-degree) of any vertex in $G$. For edge creation, zero-weight edges are constructed between consecutive time-stamped nodes and from arrival to departure nodes, which requires $\mathcal{O}(M)$ time. Additionally, each original temporal edge induces a weighted edge in $\tilde{G}$. Placing these edges in $\tilde{G}$ also requires $\mathcal{O}(M\log(d_{max}))$ time. Therefore, the overall time complexity of the graph transformation is $\mathcal{O}(M\log(d_{\max}))$.

### v2

@@
 The quantity $(t_{\ell}+\lambda_{\ell})$ is known as the ending time of $P$, denoted by $end(P)$, while $t_{1}$ is the starting time of $P$, $start(P)$.
-The duration of $P$ is defined as $dura(P)=end(P)-start(P)$.
+The duration of $P$ is defined as dura(P)=end(P)start(P)dura(P)=end(P)-start(P).
 Finally, the sum of the traversal time of each temporal edge of $P$ is known as the distance of the path, denoted as $dist(P)$.
@@
 Definition 2.2 (Dominated paths).
-Let $P$ be a temporal $u-v$ path.
-We say that $P$ is dominated if there exists another temporal $u-v$ path $P^{\prime}$ such that $\text{start}(P^{\prime})\geq\text{start}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
-We say that $P$ is distance-wise dominated if there exists a temporal $u-v$ path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
+Let $P$ be a temporal uvu-v path.
+We say that $P$ is dominated if there exists another temporal uvu-v path $P^{\prime}$ such that $\text{start}(P^{\prime})\geq\text{start}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
+We say that $P$ is distance-wise dominated if there exists a temporal uvu-v path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
 In some of our algorithms we will use a subroutine that computes all [^3] If there are paths $P,P^{\prime}$ with $start(P)=start(P^{\prime})$ and $end(P)=end(P^{\prime})$, the algorithm only stores one of them, since they can be used interchangeably for any objective.
@@
 Algorithm 1 Non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
-1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
-2:For $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
+1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
+2:For $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
 Algorithm 1 runs in time $\mathcal{O}(n+M\log(c))$, where $c=\min\{d_{in},d_{out}\}$.
@@
 Algorithm 2 Dist_non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
-1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
-2:For $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
+1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
+2:For $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
 Algorithm 2 runs in time $\mathcal{O}(n+M\log(d_{in}))$. Observe that $d_{in}$ is a bound for the amount of distance-wise non-dominated pairs $(d,a)$ (distance, arrival) between any two vertices.
@@
 We define $T=\max\limits_{b\in B}\{|T_{b}|\}$.
-Given a temporal graph $G$, vertices $x,y\in V$ and a time interval $[t_{\alpha},t_{\omega}]$, we denote the set of temporal beer paths between two vertices $x$ and $y$ starting after $t_{\alpha}$ and ending before $t_{\omega}$ as $P_{B}(x,y,[t_{\alpha},t_{\omega}])=$$\{P=\langle x=v_{1},v_{2},\dots,v_{\ell}=y\rangle:P$ is a temporal path from $x$ to $y$ such that $start(P)\geq t_{\alpha},\ end(P)\leq t_{\omega},\ \exists\ (i\in[\ell],t\in T_{v_{i}}):(v_{i}\in B,\ t\in[t_{i}+\lambda_{i},t_{i+1}])\}$.
+Given a temporal graph $G$, vertices $x,y\in V$ and a time interval $t_{\alpha},t_{\omega}]$, we denote the set of temporal beer paths between two vertices $x$ and $y$ starting after $t_{\alpha}$ and ending before $t_{\omega}$ as $P_{B}(x,y,[t_{\alpha},t_{\omega}])=$ $\{P=\langle x=v_{1},v_{2},\dots,v_{\ell}=y\rangle:P$ is a temporal path from $x$ to $y$ such that $start(P)\geq t_{\alpha},\ end(P)\leq t_{\omega},\ \exists\ (i\in[\ell],t\in T_{v_{i}}):(v_{i}\in B,\ t\in[t_{i}+\lambda_{i},t_{i+1}])\}$.
 Note that a beer path must traverse at least one beer vertex at a time in which it is active.
@@
 A temporal $x$-$y$ path $P\in{P_{B}}(x,y,[t_{\alpha},t_{\omega}])$ is a shortest beer path if $dist(P)=\min\{dist(P^{\prime}):P^{\prime}\in P_{B}(x,y,[t_{\alpha},t_{\omega}])\}$.
-Figure 1: Example in which the $x-y$ EABP/LDBP/FBP/SBP use $b_{1},\ldots,b_{4}$ respectively.
+Figure 1: Example in which the xyx-y EABP/LDBP/FBP/SBP use $b_{1},\ldots,b_{4}$ respectively.
 Numbers in parentheses denote $(t,\lambda)$ of edges and numbers in brackets denote beer vertex active times.
@@
 Definition 3.1 (MSEAP).
-Given a temporal graph $G=(V,E)$ and initial times $init[v]\in\mathbb{R}^{+}\cup\{\infty\}$, $v\in V$, we say that a temporal path $x-y$ starting from $x$ no earlier than $init[x]$ is a multiple-source earliest-arrival path to $y$ if its arrival time at $y$ is minimum among all temporal paths $v-y$, $v\in V$, starting from $v$ no earlier than $init[v]$.
+Given a temporal graph $G=(V,E)$ and initial times $init[v]\in\mathbb{R}^{+}\cup\{\infty\}$, $v\in V$, we say that a temporal path xyx-y starting from $x$ no earlier than $init[x]$ is a multiple-source earliest-arrival path to $y$ if its arrival time at $y$ is minimum among all temporal paths vyv-y, $v\in V$, starting from $v$ no earlier than $init[v]$.
 We call this arrival time the multiple-source earliest-arrival path time of $y$.
@@
 We first show that a generalization of the earliest-arrival path algorithm of [20] can compute MSEAPs.
-Algorithm 3 MSEAP$(G,[t_{\alpha},t_{\omega}],init)$
-1:A temporal graph $G$ in its edge stream representation, a time window $[t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
-2:The MSEAP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
-3:$\tau[v]\leftarrow init[v]$ for all $v\in V$
+[…]
+1:A temporal graph $G$ in its edge stream representation, a time window $t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
+2:The MSEAP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
+[…]
 4:for $e=(u,v,t,\lambda)$ in the edge stream do
@@
 Lemma 3.2 (Wu et al. [20]).
-For all nodes $x,y\in V$, if there exists some temporal $x-y$ path, then there exists some earliest-arrival $x-y$ path $\langle v_{0}=x,v_{1},v_{2},\ldots,v_{k}=y\rangle$ for which every prefix-subpath $\langle x,v_{1},v_{2},\ldots,v_{i}\rangle$ is an earliest-arrival $x-v_{i}$ path.
+For all nodes $x,y\in V$, if there exists some temporal xyx-y path, then there exists some earliest-arrival xyx-y path $\langle v_{0}=x,v_{1},v_{2},\ldots,v_{k}=y\rangle$ for which every prefix-subpath $\langle x,v_{1},v_{2},\ldots,v_{i}\rangle$ is an earliest-arrival $x-v_{i}$ path.
 {observation}
-In temporal graphs, the property described in Lemma 3.2 only holds for some $x-y$ earliest-arrival path, in contrast to classical graphs, where a similar property holds for all $x-y$ shortest paths.
+In temporal graphs, the property described in Lemma 3.2 only holds for some xyx-y earliest-arrival path, in contrast to classical graphs, where a similar property holds for all xyx-y shortest paths.
 The following lemma is analogous to Lemma 3.2 for MSEAPs.
@@
 Lemma 3.5.
-Algorithm 3 computes the MSEAP times to each vertex, within time window $[t_{\alpha},t_{\omega}]$.
-If time $\infty$ is returned for some vertex, then that vertex is unreachable from all $v\in V$ (starting no earlier than $init[v]$) within $[t_{\alpha},t_{\omega}]$.
+Algorithm 3 computes the MSEAP times to each vertex, within time window $t_{\alpha},t_{\omega}]$.
+If time $\infty$ is returned for some vertex, then that vertex is unreachable from all $v\in V$ (starting no earlier than $init[v]$) within $t_{\alpha},t_{\omega}]$.
 Proof 3.6.
-For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time (within $[t_{\alpha},t_{\omega}]$).
+For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time (within $t_{\alpha},t_{\omega}]$).
 Since $\tau[v]$ is initialized to $init[v]$, any update done by the algorithm calculates the arrival time of a valid temporal path, starting from some node $v$ at time $init[v]$.
@@
 From the above we infer that $\tau[v]=\infty$ can be returned for some $v\in V$ only if no MSEAP to $v$ exists.
-This implies that $v$ is unreachable from all $u\in V$ (starting no earlier than $init[u]$) within $[t_{\alpha},t_{\omega}]$.
+This implies that $v$ is unreachable from all $u\in V$ (starting no earlier than $init[u]$) within $t_{\alpha},t_{\omega}]$.
 We now present an EABP algorithm, which calls Algorithm 3 twice, calculating earliest-arrival path times in the first phase and EABP times in the second phase.
-Algorithm 4 EABP$(G,x,[t_{\alpha},t_{\omega}])$
-1:A temporal graph $G$, source vertex $x$, time window $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
-2:The EABP time from $x$ to every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
-3:$init[x]\leftarrow t_{\alpha},\,init[v]\leftarrow\infty$ for all $v\in V\setminus\{x\}$
-4:$\tau[v]\leftarrow\texttt{MSEAP($G,[t_{\alpha},t_{\omega}],init$)}$, for $v\in V$
-5:$init^{\prime}[v]\leftarrow\infty$ for all $v\in V$
+[…]
+1:A temporal graph $G$, source vertex $x$, time window $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
+2:The EABP time from $x$ to every vertex $v\in V$ within $t_{\alpha},t_{\omega}]$.
+[…]
+[…]
+[…]
 6:for all $b\in B$ do
-7:     $beer\_time\leftarrow\min\{time\in T_{b}\mid time\geq\tau[b]\}$triangleright $\infty$ if it does not exist
+7:  $beer\_time\leftarrow\min\{time\in T_{b}\mid time\geq\tau[b]\}$ $\triangleright$ $\infty$ if it does not exist
 8:  if $beer\_time\leq t_{\omega}$ then
 9:   $init^{\prime}[b]\leftarrow beer\_time$
-10:$\tau^{\prime}[v]\leftarrow\texttt{MSEAP($G,[t_{\alpha},t_{\omega}],init^{\prime}$)}$, for $v\in V$
+[…]
 11:return $\tau^{\prime}[v]$ for each $v\in V$
@@
 Algorithm 4 scans all edges twice. A binary search is done for each of the $k$ lists $T_{b}$. Hence, Algorithm 4 runs in time $\mathcal{O}(n+M+k\log(T))$.
-By Observation 3.1.1, the first call to Algorithm 3 correctly computes the earliest-arrival path times from $x$ to the beer vertices (within $[t_{\alpha},t_{\omega}]$).
+By Observation 3.1.1, the first call to Algorithm 3 correctly computes the earliest-arrival path times from $x$ to the beer vertices (within $t_{\alpha},t_{\omega}]$).
 Thus, the binary search on $T_{b}$ computes for each $b\in B$ the earliest departure time from $b$ s.t.
-$b$ is used as the beer vertex of a beer path from $x$ to any other node (within $[t_{\alpha},t_{\omega}]$).
+$b$ is used as the beer vertex of a beer path from $x$ to any other node (within $t_{\alpha},t_{\omega}]$).
 Denote this time as $t[b]$, $b\in B$.
@@
 This coincides with the MSEAP time to $v$ for $init[b]=t[b]$, $b\in B$, and $init[u]=\infty$, $u\in V\setminus B$.
-Hence, the second call to Algorithm 3 correctly computes the EABP times from $x$ to all nodes (within $[t_{\alpha},t_{\omega}]$), by Lemma 3.5.
+Hence, the second call to Algorithm 3 correctly computes the EABP times from $x$ to all nodes (within $t_{\alpha},t_{\omega}]$), by Lemma 3.5.
 3.1.2 A faster earliest-arrival beer path algorithm with adjacency list representation
@@
 The algorithm is similar-in-spirit to Dijkstra’s algorithm, using a min-priority queue for the nodes and their arrival times.
-Algorithm 5 MSEAP_alternative$(G,[t_{\alpha},t_{\omega}],init)$
-1:A temporal graph $G$ in adjacency list representation, a time window $[t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
-2:The MSEAP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
-3:$\tau[v]\leftarrow init[v]$ for all $v\in V$
-4:$PQ\leftarrow\emptyset$triangleright a min-priority queue
-5:$PQ.enqueue(\tau[v],v)$ for all $v\in V$triangleright Add $v$ to priority queue with $\tau[v]$ priority
+[…]
+1:A temporal graph $G$ in adjacency list representation, a time window $t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
+2:The MSEAP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
+[…]
+[…]
+[…]
 6:while $PQ\neq\emptyset$ do
@@
 Definition 3.10 (MTLDP).
-Given a temporal graph $G=(V,E)$ and final times $fin[v]\in\mathbb{R}^{+}\cup\{-\infty\}$, $v\in V$, we say that a temporal $x-y$ path arriving at $y$ no later than $fin[y]$ is a multiple-target latest-departure path from $x$ if its starting time from $x$ is maximum among all temporal paths $x-v$, $v\in V$, arriving at $v$ no later than $fin[v]$.
+Given a temporal graph $G=(V,E)$ and final times $fin[v]\in\mathbb{R}^{+}\cup\{-\infty\}$, $v\in V$, we say that a temporal xyx-y path arriving at $y$ no later than $fin[y]$ is a multiple-target latest-departure path from $x$ if its starting time from $x$ is maximum among all temporal paths xvx-v, $v\in V$, arriving at $v$ no later than $fin[v]$.
 We call this starting time the multiple-target latest-departure path time of $x$.
-The main algorithm (analogous to Algorithm 4) sets $fin[y]$ to $t_{\omega}$ and $fin[v]$ to $\infty$- ∞ for all other nodes.
+The main algorithm (analogous to Algorithm 4) sets $fin[y]$ to $t_{\omega}$ and $fin[v]$ to -$\infty$ for all other nodes.
 A MTLDP subroutine is used to calculate latest-departure times in the first phase and LDBP times in the second phase.
@@
 Lemma 3.11.
-If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is an $x-y$ FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
+If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is an xyx-y FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
 The proof of Lemma 3.11 is deferred to Appendix B.
@@
 Algorithm 1 is then used for the new graph, calculating eligible non-dominated paths from the dummy vertex to all vertices.
-Algorithm 6 FBP$(G,x,[t_{\alpha},t_{\omega}])$
-1:A temporal graph $G$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
-2:The FBP duration from $x$ to every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
-3:$L_{b}\leftarrow\texttt{Non-dom\textunderscore paths($G,x,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright calculate non-dom paths with Alg 1
+[…]
+1:A temporal graph $G$ in edge stream representation, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
+2:The FBP duration from $x$ to every vertex $v\in V$ within $t_{\alpha},t_{\omega}]$.
+[…]
 4:for $b\in B$ do
 5:  for $(s,a)\in L_{b}$ do
-6:         $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$triangleright $\infty$ if it does not exist
+6:   $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$ $\triangleright$ $\infty$ if it does not exist
 7:   if $a^{\prime}\leq t_{\omega}$ then
@@
 9:   else
-10:              Remove $(s,a)$ from $L_{b}$triangleright impossible to get beer
+10:     Remove $(s,a)$ from $L_{b}$ $\triangleright$ impossible to get beer
 11:Add dummy node $\epsilon$ to the graph. Add an edge $(\epsilon,b,s,a-s)$ for each $(s,a)\in L_{b}$ for each $b\in B$; Let $G^{\prime}=(V^{\prime},E^{\prime})$ be the modified graph
-12:$L_{v}^{\prime}\leftarrow\texttt{Non-dom\textunderscore paths($G^{\prime},\epsilon,[t_{\alpha},t_{\omega}]$)}$, for $v\in V$
+[…]
 13:for $v\in V$ do
@@
 Definition 3.12 (Inverse-distance-wise dominated path).
-Let $P$ be a temporal $u-v$ path.
-We say that $P$ is inverse-distance-wise dominated if there exists another temporal $u-v$ path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{start}(P^{\prime})\geq\text{start}(P)$, with at least one of the inequalities being strict.
+Let $P$ be a temporal uvu-v path.
+We say that $P$ is inverse-distance-wise dominated if there exists another temporal uvu-v path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{start}(P^{\prime})\geq\text{start}(P)$, with at least one of the inequalities being strict.
 Lemma 3.13.
-If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is a $x-y$ SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t.
-$\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$).
+If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is a xyx-y SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t.
+$\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$).
 The proof of Lemma 3.13 is deferred to Appendix B.
@@
 Algorithm 7 Inv_dist_non-dom_paths($G,y,[t_{\alpha},t_{\omega}]$)
-1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$.
-2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $[t_{\alpha},t_{\omega}]$.
+1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $t_{\alpha},t_{\omega}]$.
+2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $t_{\alpha},t_{\omega}]$.
 Algorithm 7 runs in time $\mathcal{O}(n+M\log(d_{out}))$.
@@
 Theorem 3.14.
-Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $[t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
+Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
 The proof of Theorem 3.14 is deferred to Appendix B.
@@
 Unlike most of our algorithms, Algorithm 8 only calculates a path between two nodes; this happens because it uses two different domination criteria (see Lemma 3.13), with each one imposing different restrictions via its respective algorithm.
-Algorithm 8 SBP$(G,x,y,[t_{\alpha},t_{\omega}])$
-1:A temporal graph $G=(V,E)$ in edge stream representation (both regular and reverse), source vertex $x$, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
-2:The SBP distance from $x$ to $y$ within $[t_{\alpha},t_{\omega}]$.
-3:$dist\leftarrow\infty$
-4:$L_{b}\leftarrow\texttt{Dist\_Non-dom\_paths($G,x,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright Alg 2 from $x$ to $b\in B$
-5:$I_{b}\leftarrow\texttt{Inv\textunderscore dist\textunderscore Non-dom\textunderscore paths($G,y,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright Alg 7 from $b\in B$ to $y$
+[…]
+1:A temporal graph $G=(V,E)$ in edge stream representation (both regular and reverse), source vertex $x$, target vertex $y$, time interval $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
+2:The SBP distance from $x$ to $y$ within $t_{\alpha},t_{\omega}]$.
+[…]
+[…]
+[…]
 6:for $b\in B$ do
 7:  for $(d,a)\in L_{b}$ do
-8:         $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$triangleright $\infty$ if it does not exist
+8:   $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$ $\triangleright$ $\infty$ if it does not exist
 9:   if $a^{\prime}\leq t_{\omega}$ then
@@
 11:   else
-12:              Remove $(d,a)$ from $L_{b}$triangleright impossible to get beer
+12:     Remove $(d,a)$ from $L_{b}$ $\triangleright$ impossible to get beer
 13:  for $(d,a)\in L_{b}$ do
-14:         Find $(d^{\prime},s)$ in $I_{b}$ with minimum $s$ such that $s\geq a$triangleright find best fitting path
+14:   Find $(d^{\prime},s)$ in $I_{b}$ with minimum $s$ such that $s\geq a$ $\triangleright$ find best fitting path
 15:   $dist\leftarrow\min\{d+d^{\prime},dist\}$
@@
 A natural application, for example, would be to quickly recompute a beer path assuming some beer shops just closed (or some new ones opened), without computing everything from scratch.
-We assume that each query contains as input a source node $x$, a target node $y$, a time interval $[t_{\alpha},t_{\omega}]$ and a Boolean array $A=[a_{1},\ldots,a_{k}]$, with $a_{i}=\texttt{true}$ iff beer vertex $b_{i}$ is active.
+We assume that each query contains as input a source node $x$, a target node $y$, a time interval $t_{\alpha},t_{\omega}]$ and a Boolean array $A=[a_{1},\ldots,a_{k}]$, with $a_{i}=\texttt{true}$ iff beer vertex $b_{i}$ is active.
 For simplicity, we assume that each beer vertex $b\in B$ is either active or inactive for the entire duration of a query.
@@
 Figure 2: Graph transformation, from $G$ in (a) to $\tilde{G}$ in (b).
-In this subsection we adopt the graph transformation approach proposed in [20] as preprocessing.
-This technique converts a temporal graph $G=(V,E)$ into a static, weighted, directed acyclic graph $\tilde{G}=(\tilde{V},\tilde{E})$ that explicitly encodes temporal constraints.
-The key idea is to model the timing of events in the temporal graph by creating time-stamped copies of each node and connecting them with appropriately weighted edges.
-The transformed graph $\tilde{G}$ contains $\mathcal{O}(M)$ vertices and edges and can be produced from $G$ in $\mathcal{O}(M\log(d_{max}))$ time.
-For each $v\in V$, the transformation creates two sets of time-stamped nodes: $\tilde{V}_{in}(v)$, which includes a node $(v,t)$ for every unique time $t$ at which an edge arrives at $v$, and $\tilde{V}_{out}(v)$, which includes a node $(v,t)$ for every time $t$ at which an edge departs from $v$.
-These sets are sorted in descending time order.
-Figures 2(a),(b) show a temporal graph $G$ and its transformed graph $\tilde{G}$, respectively.
-We defer the reader to Appendix C for more details regarding this transformation.
+[…]
 4.2.1 Fastest beer path on transformed graph
@@
 We will modify this technique to compute FBPs on $\tilde{G}$.
-For each vertex $(v,t)\in\tilde{G}$, we keep track of a $3$-valued flag, indicating whether $(v,t)$ has been visited by a beer path (value $2$) or a regular path (value $1$) or not at all (value 00), with higher values superseding smaller ones.
+For each vertex $(v,t)\in\tilde{G}$, we keep track of a $3$-valued flag, indicating whether $(v,t)$ has been visited by a beer path (value $2$) or a regular path (value $1$) or not at all (value 0), with higher values superseding smaller ones.
 We run BFSs from each $(x,t)\in S$ in descending order of $t$ (as described above), setting the flag of $(x,t)$ to $1$.
@@
 Algorithm 1 Non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
-1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
-2:For each $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
+1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
+2:For each $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
 3:for $v\in V$ do
@@
 Algorithm 2 Dist_non-dom_paths($G,x,[t_{\alpha},t_{\omega}]$)
-1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
-2:For each $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
+1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
+2:For each $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
 3:for $v\in V$ do
@@
 Algorithm 7 Inv_dist_non-dom_paths($G,y,[t_{\alpha},t_{\omega}]$)
-1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$.
-2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $[t_{\alpha},t_{\omega}]$.
+1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $t_{\alpha},t_{\omega}]$.
+2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $t_{\alpha},t_{\omega}]$.
 3:for $v\in V$ do
@@
 A.3 Omitted MTLDP and latest-departure beer path algorithms
-Algorithm 9 MTLDP$(G,[t_{\alpha},t_{\omega}],fin)$
-1:A temporal graph $G$ in its reverse edge stream representation, a time window $[t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=-\infty$, for all $v\in V$.
-2:The MTLDP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
-3:$\tau[v]\leftarrow fin[v]$ for all $v\in V$
+[…]
+1:A temporal graph $G$ in its reverse edge stream representation, a time window $t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=$-$\infty$, for all $v\in V$.
+2:The MTLDP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
+[…]
 4:for $e=(u,v,t,\lambda)$ in the reverse edge stream do
@@
 7:return $\tau[v]$ for each $v\in V$
-Algorithm 10 MTLDP_alternative$(G,[t_{\alpha},t_{\omega}],fin)$
-1:A temporal graph $G$ in adjacency list representation, a time window $[t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=-\infty$, for all $v\in V$.
-2:The MTLDP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
-3:$\tau[v]\leftarrow fin[v]$ for all $v\in V$
-4:$PQ\leftarrow\emptyset$triangleright a max-priority queue
-5:$PQ.enqueue(\tau[v],v)$ for all $v\in V$triangleright Add $v$ to priority queue with $\tau[v]$ priority
+[…]
+1:A temporal graph $G$ in adjacency list representation, a time window $t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=$-$\infty$, for all $v\in V$.
+2:The MTLDP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
+[…]
+[…]
+[…]
 6:while $PQ\neq\emptyset$ do
@@
 8:  for $u\in N_{in}(v)$ do
-9:         Let $(u,v,t,\lambda)\in E(u,v)$ be a temporal edge with maximum $t$ s.t. $t+\lambda\leq\tau[v]$triangleright binary search on $t+\lambda$ of $e\in E(u,v)$
+9:   Let $(u,v,t,\lambda)\in E(u,v)$ be a temporal edge with maximum $t$ s.t. $t+\lambda\leq\tau[v]$ $\triangleright$ binary search on $t+\lambda$ of $e\in E(u,v)$
 10:   if $t\geq t_{\alpha}$ and $t>\tau[u]$ then
@@
 13:return $\tau[v]$ for each $v\in V$
-Algorithm 11 LDBP$(G,x,[t_{\alpha},t_{\omega}])$
-1:A temporal graph $G$, source vertex $x$, a time window $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
-2:The LDBP time to $y$ from every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
-3:$fin[y]\leftarrow t_{\omega},\,init[v]\leftarrow-\infty$ for all $v\in V\setminus\{x\}$
-4:$\tau[v]\leftarrow\texttt{MTLDP($G,[t_{\alpha},t_{\omega}],fin$)}$, for $v\in V$
-5:$fin^{\prime}[v]\leftarrow-\infty$ for all $v\in V$
+[…]
+1:A temporal graph $G$, source vertex $x$, a time window $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
+2:The LDBP time to $y$ from every vertex $v\in V$ within $t_{\alpha},t_{\omega}]$.
+[…]
+[…]
+[…]
 6:for all $b\in B$ do
-7:     $beer\_time\leftarrow\max\{time\in T_{b}\mid time\leq\tau[b]\}$triangleright $\infty$- ∞ if it does not exist
+7:  $beer\_time\leftarrow\max\{time\in T_{b}\mid time\leq\tau[b]\}$ $\triangleright$ -$\infty$ if it does not exist
 8:  if $beer\_time\geq t_{\alpha}$ then
 9:   $fin^{\prime}[b]\leftarrow beer\_time$
-10:$\tau^{\prime}[v]\leftarrow\texttt{MTLDP($G,[t_{\alpha},t_{\omega}],fin^{\prime}$)}$, for $v\in V$
+[…]
 11:return $\tau^{\prime}[v]$ for each $v\in V$
@@
 We only describe the induction briefly, since it is similar to the proof of correctness of Dijkstra’s classic algorithm.
-For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time within $[t_{\alpha},t_{\omega}]$.
+For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time within $t_{\alpha},t_{\omega}]$.
 By definition of MSEAP, for the node with minimum $init[u]$ it is $\tau^{*}[u]=init[u]$; thus, the first node extracted from the queue has $\tau[u]=\tau^{*}[u]$.
-For each $v\in N_{out}(u)$, $\tau[v]$ is then updated through the $u-v$ edge that arrives at $v$ at the earliest possible time.
+For each $v\in N_{out}(u)$, $\tau[v]$ is then updated through the uvu-v edge that arrives at $v$ at the earliest possible time.
 Assuming the graph has no dominated edges, this is an edge with minimum $t$ s.t.
@@
 Lemma B.16.
-If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is an $x-y$ FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
+If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is an xyx-y FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
 Proof B.17.
-Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b,\ldots,u_{p^{\prime}}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$.
+Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b,\ldots,u_{p^{\prime}}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$.
 Define the subpaths of $P$: $Q=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b\rangle$, $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\ldots,u_{p^{\prime}}=y\rangle$.
 If $Q$ is a dominated path, then there is some other path $Q^{\prime}=\langle w_{0}=x,w_{1},\ldots,w_{f}=b\rangle$ such that $start(Q^{\prime})\geq start(Q)$ and $end(Q^{\prime})\leq end(Q)$, with at least one of the inequalities being strict.
-Concatenate $Q^{\prime}$ with $R$ to obtain a $x-y$ path $P^{\prime}$.
+Concatenate $Q^{\prime}$ with $R$ to obtain a xyx-y path $P^{\prime}$.
 By the first inequality we obtain $dura(P^{\prime})\leq dura(P)$.
 By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and that $P^{\prime}$ is a valid temporal path.
-Thus, $P^{\prime}$ is a beer path with smaller or equal duration than $P$ and is also within $[t_{\alpha},t_{\omega}]$.
+Thus, $P^{\prime}$ is a beer path with smaller or equal duration than $P$ and is also within $t_{\alpha},t_{\omega}]$.
 Applying a similar argument for $R$ completes the proof.
 Lemma B.18.
-If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is a $x-y$ SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t.
-$\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$).
+If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is a xyx-y SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t.
+$\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$).
 Proof B.19.
-Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b,\dots,u_{p^{\prime}}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at some time $t\in T_{b}$.
+Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b,\dots,u_{p^{\prime}}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at some time $t\in T_{b}$.
 Define the sub-paths of $P$: $Q=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b\rangle$ and $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\dots,u_{p^{\prime}}=y\rangle$.
-If $Q$ is distance-wise dominated (within $[t_{\alpha},t_{\omega}]$), there exists a path $Q^{\prime}=\langle w_{0}=x,w_{1},\dots,w_{f}=b\rangle$ (within $[t_{\alpha},t_{\omega}]$), such that $\text{dist}(Q^{\prime})\leq\text{dist}(Q)$ and $\text{end}(Q^{\prime})\leq\text{end}(Q)$, with at least one of the two inequalities strict.
+If $Q$ is distance-wise dominated (within $t_{\alpha},t_{\omega}]$), there exists a path $Q^{\prime}=\langle w_{0}=x,w_{1},\dots,w_{f}=b\rangle$ (within $t_{\alpha},t_{\omega}]$), such that $\text{dist}(Q^{\prime})\leq\text{dist}(Q)$ and $\text{end}(Q^{\prime})\leq\text{end}(Q)$, with at least one of the two inequalities strict.
 Concatenate $Q^{\prime}$ with $R$ to obtain a path $P^{\prime}$ from $x$ to $y$.
@@
 By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and is a valid temporal path.
-Hence, $P^{\prime}$ is a beer path within $[t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P$.
-If $R$ is inverse-distance-wise dominated (within $[t_{\alpha},t_{\omega}]$), there is a path $R^{\prime}=\langle z_{0}=b,z_{1},\dots,z_{g}=y\rangle$ (within $[t_{\alpha},t_{\omega}]$) such that $\text{dist}(R^{\prime})\leq\text{dist}(R)$ and $\text{start}(R^{\prime})\geq\text{start}(R)$, again with at least one inequality being strict.
-Concatenate $Q^{\prime}$ with $R^{\prime}$ to obtain a $x-y$ path $P^{\prime\prime}$.
+Hence, $P^{\prime}$ is a beer path within $t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P$.
+If $R$ is inverse-distance-wise dominated (within $t_{\alpha},t_{\omega}]$), there is a path $R^{\prime}=\langle z_{0}=b,z_{1},\dots,z_{g}=y\rangle$ (within $t_{\alpha},t_{\omega}]$) such that $\text{dist}(R^{\prime})\leq\text{dist}(R)$ and $\text{start}(R^{\prime})\geq\text{start}(R)$, again with at least one inequality being strict.
+Concatenate $Q^{\prime}$ with $R^{\prime}$ to obtain a xyx-y path $P^{\prime\prime}$.
 By the first inequality we obtain $\text{dist}(P^{\prime\prime})\leq\text{dist}(P^{\prime})$.
 By the second inequality, we obtain that $P^{\prime\prime}$ is a valid temporal beer path.
-Therefore, $P^{\prime\prime}$ is a beer path within $[t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P^{\prime}$.
+Therefore, $P^{\prime\prime}$ is a beer path within $t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P^{\prime}$.
 Theorem B.20.
-Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $[t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
+Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
 Proof B.21.
-Let $P=\langle v_{0}=v,v_{1},\dots,v_{p}=y\rangle$ be an inverse-distance-wise non-dominated path from $v$ to $y$ within $[t_{\alpha},t_{\omega}]$.
+Let $P=\langle v_{0}=v,v_{1},\dots,v_{p}=y\rangle$ be an inverse-distance-wise non-dominated path from $v$ to $y$ within $t_{\alpha},t_{\omega}]$.
 For $i\in\{0,\ldots,p-1\}$, let $s_{i}$ be the time in which $P$ departs from $v_{i}$ and $d_{i}$ be the distance between $v_{i}$ and $y$ in $P$.
@@
 From the above induction we infer that $I_{v}$ contains an element $(d,s)$ with $d\leq d_{0}$ and $s\geq s_{0}$.
-By assumption, $P$ is an inverse-distance-wise non-dominated path within $[t_{\alpha},t_{\omega}]$, which implies that neither of the above inequalities can be strict.
+By assumption, $P$ is an inverse-distance-wise non-dominated path within $t_{\alpha},t_{\omega}]$, which implies that neither of the above inequalities can be strict.
 Hence, $I_{v}$ contains $(d_{0},s_{0})$, which finishes the correctness proof.
@@
 Appendix C The graph transformation of Wu et al.
-In this section we describe in detail the graph transformation of [20], which converts a temporal graph $G=(V,E)$ into an equivalent static, weighted, directed acyclic graph $\tilde{G}=(\tilde{V},\tilde{E})$.
+[…]
 For each vertex $v\in V$, the transformation creates two sets of time-stamped nodes: $\tilde{V}_{in}(v)$, which includes a node $(v,t)$ for every unique time $t$ at which an edge arrives at $v$, and $\tilde{V}_{out}(v)$, which includes a node $(v,t)$ for every time $t$ at which an edge departs from $v$.